In a random sample of 60 students in adult education classes on a large college campus, it was found that 75% of the sample members were not in the market for cigarettes. what would be the probability of observing a sample proportion that is less than this if in fact 70% of all adult education students on campus are not in the market for cigarettes?
Answer #1

Solution

=  [p ( 1 - p ) / n] =   [(0.75 * 0.25) / 60 ] = 0.05590

P( < 0.70) =

= P[( - ) / < (0.70 - 0.75) / 0.0559]

= P(z < 0.89)

= 0.8133